Help me solve? Don’t think I got the right answer
The table lists the smoking habits of a group of college students. Sex Non-smoker Regular Smoker Heavy Smoker Total 5 186 12 220 17 406 135 187 322 46 21 67 Man Woman Total If a student is chosen at random, find the probability of getting someone who is a man or a non-smoker Round your answer to three decimal places O A. 0.942 OB, 0.835 o c. 0.947 ra D. 0.919 Click to select your answer 0 20 F3
Answer #1

Option D

Probability = Favourable Outcomes (F) / Total Outcomes (T)

Total Outcomes = 406C1 = 406

Favourable Outcomes will be Either a Man or a Non Smoker = Man + Non Smoker – Man and Non Smoker = 186C1 + 322C1 - 135C1 = 186+322-135 = 373

Hence, Probability = 373/406 = 0.9187 ~ 0.919